From GlueXWiki
Jump to: navigation, search

November 8, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • First package testing (Vlad)
  2. Corrosion tests, corroded solder on Viton sample (Vlad)
  3. Engineering (Bill)
    • Bubbler issue and pressure control [1]
  4. Electronics (Chris, Nick)
  5. Third package tests FDC E-log (Beni, Simon)
    • Moving to Blue Crab?
  6. Other


Participants: Bill, Dave, Chris, Nick, Vlad, Simon, Beni, and Lubomir.


- Casey and Tina strung wire frame #30 on Saturday. Dave put the Kapton dots and did position measurements. Casey will come again this Saturday to glue the wires. Anatoly is preparing the old cathode frames for package #2 and will finish soon.

- Vlad is testing package #1. Currents on cell #4 300nA, #3 350nA, #2 170nA, #1 (just started) 500nA. For comparison 3 weeks ago we had 900nA on cell #4. All the channels so far are working fine. Vlad found two pre-amps oscillating.

Cathode corrosion test

- The testing chamber was opened. Sample #1 EPDM failed after 12 days, Sample #2 Viton failed after 4 days, Sample #3 Viton+Apiezon didn't fail. After closer look at Sample #2 it turned out that the solder on one side of the trace corroded, or was somehow damaged during soldering and later at this high temperature and humidity failed. In gas EPDM failed after 8/10 days (we had two samples). So maybe it is only the EPDM itself responsible for the corrosion. The good news: no corrosion observed yet when using Viton. Also, the fail of sample #2 indicates problems when soldering on the thin copper; good we didn't do it!


- Bill has ordered the new SMC chiller; the fluorinert costs $1.7k. Working on the documentation for the pressure system.

- Over-pressure problem over the weekend described in the Elog-#95. No issues with the chamber found due to the over-pressure. The bubbler has a problem that we observed already before: at some higher pressure (for package #1 in happened at a flow of 220ccpm) it stops bubbling and the pressure in the chamber increases; it can reach very high values as it happened over the weekend. So far most likely explanation: when the bubble reaches the bottom of the bubbler it makes tiny gas channel to the surface and the bubble stops expanding. This state is stable even at very high pressure and the impedance of the bubbler is very high. We will try four other bubbler types. Bill proposed to study the phenomenon without the chamber by measuring the hysteresis curve pressure_vs_flow of the bubbler itself.


- Vlad fond two pre-amp cards oscillating, no matter in what position on the cathodes they are. The pre-amp testing set-up has small capacitance (short cable) at the input unlike the detector. Will discuss with Cody changes in the set-up to better simulate the detector.

Third package tests

- Beni: if just looking at the drift times (without converting to drift distances) and requiring the same drift time from two parallel wires, he obtained ~250 micron averaged resolution.

- External chamber information was included in the hddm file, but Simon doesn't see correlations with the FDC data, except roughly the expected position.

- Moving to Blue Crab: according to Ron Bartek (Bill) fixing the temperature in 126 using ceiling fans will cost $19k. Installing the power at Blue Crab for the testing set-up will cost $7k, but including the power for the chiller that we will have to do in 126, as well.